Lysine, an essential amino acid in the human body, contains C, H, O, and N. In one experiment, the complete combustion of 2.175 g of lysine gave 3.94 g CO2 and 1.89 g H2O. In a separate experiment, 1.873 g of lysine gave 0.436 g NH3. (a) Calculate the empirical formula of lysine. (b) The approximate molar mass of lysine is 150 g. What is the molecular formula of the compound?


1
Expert's answer
2021-10-05T01:44:04-0400

Molar mass of CO2 = 44.01

3.94/44.01= 0.09

Molar mass of H2O= 18.01528 g

1.89/18.01528= 0.1

Molar mass of NH3 = 17.031 g

0.436/17.031= 0.026

CO2H2O5NH3